OpenStudy (anonymous):

Given complex number z=3+4i Write z in polar form & Write z^5 in polar form

OpenStudy (anonymous):

you want \[r\left(\cos(\theta)+i\sin(\theta)\right)\] or \[re^{i\theta}\]?

OpenStudy (anonymous):

r(cos(theta)+isin(theta)) please

OpenStudy (amistre64):

how far is the point (3,4) from the origin? and what angle does it make with the positive x axis?

OpenStudy (anonymous):

\[r=\sqrt{a^2+b^2}\]

OpenStudy (anonymous):

r=5

OpenStudy (amistre64):

and that arctangent of 4/3 is?

OpenStudy (anonymous):

53.1. so z=5(cos53+isin53) and z^5=3125(cos265+isin265)

OpenStudy (amistre64):

seems reasonable enough to me :)